OpenStudy (anonymous):

2x - 40 = 2y 2y = 64 - 6x Using the two equations above, solve for y.

OpenStudy (kaitlyn_nicole):

so to find a variable, i was taught to substitute 0 for the other.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if you divide the second equation all the way across by 2, you get \[y=32-3x\]

OpenStudy (misty1212):

try dividing the first one by 2 and see what you get

OpenStudy (misty1212):

then set the two expressions for \(y\) equal, solve for \(x\)

OpenStudy (anonymous):

y=x-20

OpenStudy (misty1212):

right so now solve \[32-3x=x-20\] for \(x\) in three steps

OpenStudy (misty1212):

a) add \(3x\) to both sides b) add \(20\) to both sides c) divide both sides by \(4\)

OpenStudy (retireed):

I think this way is easier to visualize. Rewrite using algebra so the x and y terms line up... ax + by = c1 dx - ey = c2 in this form elimination of an x or y term is possible by addition or subtraction.

OpenStudy (anonymous):

x=13

OpenStudy (retireed):

2x - 40 = 2y ------> 2x - 2y = 40 2y = 64 - 6x -----> 6x + 2y = 64 Now if I add the two equations.... The y term will equal 0

OpenStudy (misty1212):

yes \(x=13\)

OpenStudy (misty1212):

and since \(y=x-20\) that tells you \(y=13-20\)

OpenStudy (retireed):

2x - 40 = 2y ------> 2x - 2y = 40 2y = 64 - 6x -----> 6x + 2y = 64 add 8x +0y = 104 x = 104 / 8 as seen above x = 13

OpenStudy (retireed):

2x - 40 = 2y 2y = 64 - 6x --> x = 13 --> 2y = 64 - 78 --> y = 32 - 39 y = -7 Check by plugging answers into the other equation .... 2x - 40 = 2y x =13 y = -7 26 - 40 = -14 CHECKS